Kokomo uses a standard costing system. For production of a fuel cell, standards call for 4 model RN-3 valves per cell costing $2,340 each. During 2018, the company purchased 260 valves for $642,200. The company used 252 valves in the production of 61 fuel cells (8 valves were ruined through installation error).

Calculate material price and quantity variances for the model RN-3 valves. (Enter all variances as a positive number.)

Material Price Variance $

Neither Unfavorable nor FavorableFavorableUnfavorable

Material Quantity Variance $

FavorableNeither Unfavorable nor FavorableUnfavorable
Answer #1

Materials Price Variance = (SP-AP) * AQ

= (2340 - 642,200/260) * 260

= 33,800 Unfavorable

.

Materials Quantity Variance = (SQ-AQ) * SP

= (61*4 - 252) * 2340

= 18,720 Unfavorable